OpenStudy (anonymous):

find the coordinates of a second point on the graph of a function f if the given point is on the graph and the function is (a) even and (b) odd.?? (4,9)

OpenStudy (anonymous):

f is even means \[f(-x)=f(x)\] so if \[f(4)=9\] then \[f(-4)=9\] as we;;

OpenStudy (anonymous):

*as well

jimthompson5910 (jim_thompson5910):

a) Since the function is even, we know that f(-x) = f(x) Because f(4) = 9, we can also say that f(-4) = 9 as well (since x = 4) b) If the function is odd, then f(-x) = -f(x) So if f(4) = 9, then f(-4) = -9

OpenStudy (anonymous):

thanks i c now

OpenStudy (anonymous):

Satellite 5 questions...I posted.....